# New News: System-2 Fine-tuning for Robust Integration of New Knowledge

### v2

@@ 1 Introduction
Learning new knowledge in a consistent and continuous way is one of the most important cognitive abilities. While large language models (Brown et al., 2020; OpenAI, 2024a; Team, 2024b; Dubey et al., 2024; Yang et al., 2024) have been successful in crushing knowledge and problem-solving focused benchmarks (Hendrycks et al., 2021a; b; Rein et al., 2023), these benchmarks do not measure the ability to successfully adapt one’s belief when internalizing new information, which is arguably a hallmark of general intelligence. Interestingly, current models demonstrate impressive in-context learning abilities and can process novel information efficiently when the new information is given in context (Wei et al., 2023; Team, 2024a; Lampinen et al., 2022; Park et al., 2024); yet, it remains a challenge to consolidate the knowledge in weights (Snell et al., 2022; Berglund et al., 2024; Guan et al., 2025). For example, upon learning the news that Trump won the 2024 US election, most people can instinctively update their world models, propagate implications and react accordingly. Given such news as context, language models are able to process its implication adequately as well through its chain of thoughts; however, it is hard to distill these implications back into its weights via fine-tuning (FT), as past works have shown the unreliability of fine-tuning as a knowledge injection technique (Mitchell et al., 2022a; Meng et al., 2022; Berglund et al., 2024).
@@ 2 Related works
Our work shares motivation with knowledge editing (Zhang et al., 2024; Wang et al., 2024a) in updating models’ factual understandings and world models. However, a major difference is that our goal is not to precisely edit a relational fact in the model, but rather to integrate new and non-counterfactual information appropriately into the model. In this regard, both the New News dataset and Sys2-FT protocols are different from the approaches explored in the knowledge editing literature. Traditional editing datasets and benchmarks such as zsRE (Levy et al., 2017) and Wikibio (Lebret et al., 2016; Zhang et al., 2024) tend to focus on simple subject-relation-object pair edits, whereas our evaluation dataset consists of more realistic and comprehensive updates on news; classic model editing methods such as ROME (Meng et al., 2022) and MEND (Mitchell et al., 2022b) also tend to localize the edit in the model a priori, whereas the System-2 Fine-tuning method is more natural and not constrained in any way beyond the choice of data augmentation protocols. In fact, due to the aforementioned points, knowledge editing techniques often suffer from producing incorrect ripple effects and incoherent downstream reasoning in multi-hop questions (Cohen et al., 2023; Zhong et al., 2023), and more importantly the target effect is often not even well defined (Hase et al., 2024).
@@
There is evidence that LLMs have internal world model representations (Li et al., 2022; Gurnee & Tegmark, 2023; Hazineh et al., 2023; Vafa et al., 2024) and form beliefs about the world (Hase et al., 2021; Zhu et al., 2024; Scherrer et al., 2023). While there have been attempts to revise and modify the beliefs of the model when presented with new information (Hase et al., 2021; Wilie et al., 2024), they are usually limited to simple counterfactual examples and involve architecture dependent edits Meng et al. (2022). On the other hand, our news dataset mainly consists of plausible scenarios with well-known entities and the Sys2-FT method is architecture agonistic.
@@
Another related line of works with rich literature is knowledge distillation (Buciluǎ et al., 2006; Ba & Caruana, 2014; Hinton et al., 2015) which has been studied in the field of LLMs in recent years (Xu et al., 2024; Gu et al., 2023; Wang et al., 2022; Zelikman et al., 2022; Agarwal et al., 2024). Closest to our work are context distillation works, which aim to distill the knowledge from a model given some context into its weights (Huang et al., 2022; Snell et al., 2022; Padmanabhan et al., 2023; Wang et al., 2024b; Kujanpää et al., 2024). In fact, our Sys2-FT protocols, especially implication and Self-QA can be thought of as context distillation with a simplified training procedure but a more agressive data augmentation scheme than previous works.
@@
Finally, many works explored the performance of ICL vs. FT on various tasks (Sun et al., 2023; Mosbach et al., 2023; Pecher et al., 2024; Balaguer et al., 2024; Feng et al., 2024), but findings are inconclusive as performances are often dependent on dataset, task and model. We systematically compare ICL vs. FT on our New News dataset and argue that the fine-tuning baseline can be significantly improved via Sys2-FT.
@@ 4 System-2 Fine-tuning
Humans and animals effectively consolidate new memories through deliberate rehearsal, rephrasing, and self-explanation (Diekelmann & Born, 2010; Craik & Lockhart, 1972; Slamecka & Graf, 1978; Chi et al., 1994). Neuroscience studies further reveal that offline replay (Ericsson et al., 1993) and schema integration play crucial roles in assimilating information into existing knowledge structures (Wilson & McNaughton, 1994; Tse et al., 2007; Van Genderen et al., 2012). Additionally, theoretical frameworks such as the complementary learning systems model (McClelland et al., 1995; O’Reilly et al., 2014) highlight the necessity of an integrative consolidation phase, involving deeper, controlled reprocessing, to embed knowledge robustly. Inspired by these findings on memory consolidation and knowledge integration, we introduce a general approach termed System-2 Fine-tuning (Sys2-FT), wherein models actively rehearse, paraphrase, and generate self-explanations about newly encountered information in context, the data of which is then utilized for fine-tuning, as shown in the pipeline of Fig. 1.
@@ 4.2 Results of System-2 Fine-tuning
We system-2 fine-tune the Qwen 2.5 family of models (Yang et al., 2024) on the New News dataset, and find that Sys2-FT methods consistently achieve higher performance across dataset domains and model sizes compared to naive-FT. Especially, we find our Self-QA protocol to be significantly stronger than other protocols, almost achieving near ICL performance in math and coding splits for larger models, as shown in Fig. 4. In general, we find quantitative domains (Math and Coding) to benefit more for Sys2-FT.
@@
Therefore, we confirm that rephrasing and replaying the information during fine-tuning, especially Self-QAs is crucial for models to integrate and internalize new knowledge. In fact, LLMs fail to learn the concepts that have not appeared in multiple contexts for multiple times similarly in pre-training (Kandpal et al., 2023; Allen-Zhu & Li, 2023).
@@ 6 General Capability Preservation and Continual Learning
Figure 7: General Capability Evaluation after System-2 Finetuning We evaluated Qwen-7B and 14B models before and after Self-QA Sys2-FT on MMLU-Pro (Wang et al., 2024c), GSM8K (Cobbe et al., 2021) and IFEval (Zhou et al., 2023). The general knowledge and instruction-following capabilities are preserved during training, except for GSM8K, which shows a slight degradation in the discoveries domain. See App. C.1 for more detailed results on dynamics and subdomain evaluations.
As model fine-tuned with the Sys2-FT method on our dataset achieve high evaluation performance on the new knowledge, an important question that remains is to what extent it will harm the preexisting capabilities of the model. In Fig. 7, we show that general knowledge and instruction following capabilities are well preserved after Self-QA System-2 Fine-tuning. In another set of experiments (Fig. 8), after fine-tuning on one subdomain of the New News dataset and merging with that LoRA (Hu et al., 2021) adapter, we continually fine-tune the model on another subdomain of the dataset and observe relatively small forgetting of the knowledge just learned. These experiments suggest that preexisting knowledge and capabilities are preserved after Self-QA System-2 fine-tuning with relatively small forgetting, as the model continually learns new information.
Figure 8: Recency: forgetting of previous news when training for other news on Qwen-7B. We show here the results of Sys2-FT with the Self-QA protocol on one news category, followed by another. Specifically, we explore 4 training combinations where the model is first trained with one category (eg. math) and then merged the LoRA (Hu et al., 2021) adapter, followed by training with another category (eg. coding). We evaluate the performance of both categories during continual learning of the two dataset categories. The solid lines show the category currently being trained on, and the dashed lines show the category that is not currently being trained.
@@ 7 Analysis of Sys2-FT across Data Generation models
To isolate the model size effect, we conduct Self-QA FT experiments across model scales with data generated all other model scales as well. We focus on the Self-QA protocol as it is the most stable and robust across data splits and model sizes (Fig. 4). As shown in Fig. 9, we notice that the successful models lie in a rectangular region in the heatmap, which suggests that: 1) Successful integration of new knowledge requires both a sufficiently good model and high quality data, as opposed to a smooth trade off; 2) Stronger models fine-tuned with lower-quality data generated by weak models can surpass the performance of weak model itself, a phenomenon reminiscent of weak-to-strong generalization (Burns et al., 2023). Additional details on data quality generated by models of different scales and overlap with the evaluation questions can be found in Appendix C.4.
@@ Appendix A Dataset Curation
This is an informal comment: We tried to generate news or questions using OpenAI’s o1 or o3 (OpenAI, 2024b), but found that they are largely unable to generate news for which it is easy to generate well-defined downstream questions. Often, the generated news does meet the criteria for being plausible and having downstream effects, but they are quite vague so that creating evaluation questions is challenging. There were also largely unable to generate good evaluation questions, and usually either made a question which cannot be presented standalone without the news (even thought prompted to not do so) or having hidden cues to deduce the answer without the news. The answer options were also hard to use as they often generated sets of answer where once can exploit the fact that there is only one correct answer.
@@ B.1 Models
We use the Qwen2.5 familiy of models (Yang et al., 2024) for all experiments. In specific, we use Qwen2.5-xxB-Instruct for $x\in[0.5,1.5,3,7,14,32]$.
@@ B.2 Hyperparameters
Table 2: Compute settings used for the experiments. All models are Qwen2.5-xxB-Instruct models. All GPUs used are Nvidia H100 80GB GPUs. All data parallel training runs use accelerate (Gugger et al., 2022). Note that model parallelism treats the 4 GPUs as “one device”, so the Per GPU Batch Size is simply the forward pass batch size. All generations use vLLM (Kwon et al., 2023).
@@ B.4 Fine-Tuning
The fine-tuning hyperparameters are in Tab. 1. As seen above in Sec. B.3, we generate 1024 user-assistant conversations per news. We stack all data in the same split (domain) to construct the fine-tuning dataset, and get $15\times 1024=15360$ rows (conversations). We fine-tune all models with the standard next token cross entropy loss. We only compute the loss on tokens corresponding to the assistant’s turn in the conversation. We train for 4 epochs and save 80 checkpoints throughout the runs. While this might seem like a lot of storage, since we only save the LoRA adapters (Hu et al., 2021), each checkpoint is only $80\text{MB}$ even for the biggest model (Qwen2.5-32B-Instruct).
@@ B.5 Evaluation
We evaluate every model’s accuracy using the downstream questions from the New News dataset. We evaluate each question 5 times, while randomly mixing the A,B,C,D ordering of the options to avoid bias (Pezeshkpour & Hruschka, 2023; Zheng et al., 2024).
@@ D.2 System 2 fine-tuning accuracies across model scales and data splits
We also quantify the performance gap recovered defined in Burns et al. (2023), for all models, data splits and protocols in Fig. 28.

### v3

@@ 1 Introduction
Learning new knowledge in a consistent and continuous way is one of the most important cognitive abilities. While large language models (brown2020language; openai2024gpt4technicalreport; geminiteam2024geminifamilyhighlycapable; dubey2024llama3herdmodels; yang2024qwen2) have been successful in crushing knowledge and problem-solving focused benchmarks (hendrycks2021measuringmassivemultitasklanguage; hendrycks2021measuringmathematicalproblemsolving; rein2023gpqagraduatelevelgoogleproofqa), these benchmarks do not measure the ability to successfully adapt one’s belief when internalizing new information, which is arguably a hallmark of general intelligence. Interestingly, current models demonstrate impressive in-context learning abilities and can process novel information efficiently when the new information is given in context (wei2023largerlanguagemodelsincontext; geminiteam2024gemini15unlockingmultimodal; lampinen2022languagemodelslearnexplanations; park2024iclrincontextlearningrepresentations); yet, it remains a challenge to consolidate the knowledge in weights (snell2022learningdistillingcontext; berglund2024reversalcursellmstrained; guan2025deliberativealignmentreasoningenables). For example, upon learning the news that Trump won the 2024 US election, most people can instinctively update their world models, propagate implications and react accordingly. Given such news as context, language models are able to process its implication adequately as well through its chain of thoughts; however, it is hard to distill these implications back into its weights via fine-tuning (FT), as past works have shown the unreliability of fine-tuning as a knowledge injection technique (mitchell2021fast; meng2022locating; berglund2024reversalcursellmstrained).
@@ 2 Related works
Our work shares motivation with knowledge editing (zhang2024comprehensivestudyknowledgeediting; wang2024knowledge) in updating models’ factual understandings and world models. However, a major difference is that our goal is not to precisely edit a relational fact in the model, but rather to integrate new and non-counterfactual information appropriately into the model. In this regard, both the New News dataset and Sys2-FT protocols are different from the approaches explored in the knowledge editing literature. Traditional editing datasets and benchmarks such as zsRE (levy-etal-2017-zero) and Wikibio (DBLP:journals/corr/LebretGA16; zhang2024comprehensivestudyknowledgeediting) tend to focus on simple subject-relation-object pair edits, whereas our evaluation dataset consists of more realistic and comprehensive updates on news; classic model editing methods such as ROME (meng2022locating) and MEND (mitchell2022fastmodeleditingscale) also tend to localize the edit in the model a priori, whereas the System-2 Fine-tuning method is more natural and not constrained in any way beyond the choice of data augmentation protocols. In fact, due to the aforementioned points, knowledge editing techniques often suffer from producing incorrect ripple effects and incoherent downstream reasoning in multi-hop questions (cohen2023evaluatingrippleeffectsknowledge; zhong2023mquake), and more importantly the target effect is often not even well defined (hase2024fundamentalproblemsmodelediting).
@@
There is evidence that LLMs have internal world model representations (li2022emergent; gurnee2023language; hazineh2023linear; vafa2024evaluating) and form beliefs about the world (hase2021language; zhu2024language; scherrer2023evaluating). While there have been attempts to revise and modify the beliefs of the model when presented with new information (hase2021language; wilie2024belief), they are usually limited to simple counterfactual examples and involve architecture dependent edits meng2022locating. On the other hand, our news dataset mainly consists of plausible scenarios with well-known entities and the Sys2-FT method is architecture agonistic.
@@
Another related line of works with rich literature is knowledge distillation (buciluǎ2006model; ba2014deep; hinton2015distilling) which has been studied in the field of LLMs in recent years (xu2024survey; gu2023minillm; wang2022self; zelikman2022star; agarwal2024policy). Closest to our work are context distillation works, which aim to distill the knowledge from a model given some context into its weights (huang2022incontextlearningdistillationtransferring; snell2022learningdistillingcontext; padmanabhan2023propagating; wang2024self; kujanpaa2024knowledge). In fact, our Sys2-FT protocols, especially implication and Self-QA can be thought of as context distillation with a simplified training procedure but a more agressive data augmentation scheme than previous works.
@@
Finally, many works explored the performance of ICL vs. FT on various tasks (sun2023pushing; mosbach2023few; pecher2024comparing; balaguer2024rag; feng2024extractive), but findings are inconclusive as performances are often dependent on dataset, task and model. We systematically compare ICL vs. FT on our New News dataset and argue that the fine-tuning baseline can be significantly improved via Sys2-FT.
@@ 4 System-2 Fine-tuning
Humans and animals effectively consolidate new memories through deliberate rehearsal, rephrasing, and self-explanation (diekelmann2010memory; craik1972levels; slamecka1978generation; chi1994eliciting). Neuroscience studies further reveal that offline replay (ericsson1993role) and schema integration play crucial roles in assimilating information into existing knowledge structures (wilson1994reactivation; tse2007schemas; van2012theoretical). Additionally, theoretical frameworks such as the complementary learning systems model (mcclelland1995there; o2014complementary) highlight the necessity of an integrative consolidation phase, involving deeper, controlled reprocessing, to embed knowledge robustly. Inspired by these findings on memory consolidation and knowledge integration, we introduce a general approach termed System-2 Fine-tuning (Sys2-FT), wherein models actively rehearse, paraphrase, and generate self-explanations about newly encountered information in context, the data of which is then utilized for fine-tuning, as shown in the pipeline of Fig. 1.
@@ 4.2 Results of System-2 Fine-tuning
We system-2 fine-tune the Qwen 2.5 family of models (yang2024qwen2) on the New News dataset, and find that Sys2-FT methods consistently achieve higher performance across dataset domains and model sizes compared to naive-FT. Especially, we find our Self-QA protocol to be significantly stronger than other protocols, almost achieving near ICL performance in math and coding splits for larger models, as shown in Fig. 4. In general, we find quantitative domains (Math and Coding) to benefit more for Sys2-FT.
@@
Therefore, we confirm that rephrasing and replaying the information during fine-tuning, especially Self-QAs is crucial for models to integrate and internalize new knowledge. In fact, LLMs fail to learn the concepts that have not appeared in multiple contexts for multiple times similarly in pre-training (kandpal2023large; allen2023physics).
@@ 6 General Capability Preservation and Continual Learning
Figure 7: General Capability Evaluation after System-2 Finetuning We evaluated Qwen-7B and 14B models before and after Self-QA Sys2-FT on MMLU-Pro (wang2024mmluprorobustchallengingmultitask), GSM8K (cobbe2021training) and IFEval (zhou2023instructionfollowingevaluationlargelanguage). The general knowledge and instruction-following capabilities are preserved during training, except for GSM8K, which shows a slight degradation in the discoveries domain. See App. C.1 for more detailed results on dynamics and subdomain evaluations.
As model fine-tuned with the Sys2-FT method on our dataset achieve high evaluation performance on the new knowledge, an important question that remains is to what extent it will harm the preexisting capabilities of the model. In Fig. 7, we show that general knowledge and instruction following capabilities are well preserved after Self-QA System-2 Fine-tuning. In another set of experiments (Fig. 8), after fine-tuning on one subdomain of the New News dataset and merging with that LoRA (hu2021loralowrankadaptationlarge) adapter, we continually fine-tune the model on another subdomain of the dataset and observe relatively small forgetting of the knowledge just learned. These experiments suggest that preexisting knowledge and capabilities are preserved after Self-QA System-2 fine-tuning with relatively small forgetting, as the model continually learns new information.
Figure 8: Recency: forgetting of previous news when training for other news on Qwen-7B. We show here the results of Sys2-FT with the Self-QA protocol on one news category, followed by another. Specifically, we explore 4 training combinations where the model is first trained with one category (eg. math) and then merged the LoRA (hu2021loralowrankadaptationlarge) adapter, followed by training with another category (eg. coding). We evaluate the performance of both categories during continual learning of the two dataset categories. The solid lines show the category currently being trained on, and the dashed lines show the category that is not currently being trained.
@@ 7 Analysis of Sys2-FT across Data Generation models
To isolate the model size effect, we conduct Self-QA FT experiments across model scales with data generated all other model scales as well. We focus on the Self-QA protocol as it is the most stable and robust across data splits and model sizes (Fig. 4). As shown in Fig. 9, we notice that the successful models lie in a rectangular region in the heatmap, which suggests that: 1) Successful integration of new knowledge requires both a sufficiently good model and high quality data, as opposed to a smooth trade off; 2) Stronger models fine-tuned with lower-quality data generated by weak models can surpass the performance of weak model itself, a phenomenon reminiscent of weak-to-strong generalization (burns2023weak). Additional details on data quality generated by models of different scales and overlap with the evaluation questions can be found in Appendix C.4.
@@ Appendix A Dataset Curation
This is an informal comment: We tried to generate news or questions using OpenAI’s o1 or o3 (openai2024openaio1card), but found that they are largely unable to generate news for which it is easy to generate well-defined downstream questions. Often, the generated news does meet the criteria for being plausible and having downstream effects, but they are quite vague so that creating evaluation questions is challenging. There were also largely unable to generate good evaluation questions, and usually either made a question which cannot be presented standalone without the news (even thought prompted to not do so) or having hidden cues to deduce the answer without the news. The answer options were also hard to use as they often generated sets of answer where once can exploit the fact that there is only one correct answer.
@@ B.1 Models
We use the Qwen2.5 familiy of models (yang2024qwen2) for all experiments. In specific, we use Qwen2.5-xxB-Instruct for $x\in[0.5,1.5,3,7,14,32]$.
@@ B.2 Hyperparameters
Table 2: Compute settings used for the experiments. All models are Qwen2.5-xxB-Instruct models. All GPUs used are Nvidia H100 80GB GPUs. All data parallel training runs use accelerate (accelerate). Note that model parallelism treats the 4 GPUs as “one device”, so the Per GPU Batch Size is simply the forward pass batch size. All generations use vLLM (kwon2023efficient).
@@ B.4 Fine-Tuning
The fine-tuning hyperparameters are in Tab. 1. As seen above in Sec. B.3, we generate 1024 user-assistant conversations per news. We stack all data in the same split (domain) to construct the fine-tuning dataset, and get $15\times 1024=15360$ rows (conversations). We fine-tune all models with the standard next token cross entropy loss. We only compute the loss on tokens corresponding to the assistant’s turn in the conversation. We train for 4 epochs and save 80 checkpoints throughout the runs. While this might seem like a lot of storage, since we only save the LoRA adapters (hu2021loralowrankadaptationlarge), each checkpoint is only $80\text{MB}$ even for the biggest model (Qwen2.5-32B-Instruct).
@@ B.5 Evaluation
We evaluate every model’s accuracy using the downstream questions from the New News dataset. We evaluate each question 5 times, while randomly mixing the A,B,C,D ordering of the options to avoid bias (pezeshkpour2023largelanguagemodelssensitivity; zheng2024largelanguagemodelsrobust).
@@ D.2 System 2 fine-tuning accuracies across model scales and data splits
We also quantify the performance gap recovered defined in burns2023weak, for all models, data splits and protocols in Fig. 28.